3. For each of the following matrices, a. Determine the characteristic polynomial corresponding to the matrix. b. Find the eiCould you please just solve Question (i) A: Thanks
Answer #1

/ 2 -3 : A= 11 - 2! - 3 2 1 @ the charac tenistic polynomial is - JA-113!=0 | A- #131-0 og Dahi ie (2-2 -3 na -0 1 11 -2-3 -3© Let Eo be the eigen space Corzesponding to the e. v o ie null space of the matrix Eo = K (A-0.1) =K-A we reduce the matrixSimilarly o for Ei = K (A-1.I) =k ( A-I) wall -3 17 - A-1 = -3.|| Clearly, the now neduced form is L 0 0 0 Thus, the solutionLet 20= C 22= d. 23 = -2+3d. men x. = 1 -- []*[i] whene, cand d ame Constant @ we choose, del on The condition for eigen valu